

********************************************************************
UARS WEEKLY STATUS REPORT - 26 MAY 2000
********************************************************************
***********
SOLSTICE
***********
We continue to gather enough data to meet all the SOLSTICE science
objectives during this period of reduced tape recorder operations.
SOLSTICE operates normally, collecting solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. Level 2 processing is halted due to bugs in the OA_SSPP_ATT
service. CDHF and CSC personnel are working on the problem.
***********
SUSIM
***********
Nothing received.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM AXIS measured X-rays from half of its camera (8 pixels) from
May 3 to May 6 and again beginning May 20. The full AXIS instrument measured
X-rays between May 6 and May 12. At the present time, PEM energy deposition
is not being generated due to failures in the inversion code. These are
currently being examined.
The PEM spectrometers continue to be active under PMON control*. At the
present time, PEM energy deposition is not being generated. Recent data
configuration changes have affected PEM software causing multiple failures.
These failures are being reviewed by the PEM team. In order to correct for new
telemetry requirements, PEM will need to update processing code with a patch
delivery. Updated HEPS calibration parameters were finalized and are ready for
delivery. Updated MEPS quality calibration files were finalized and are ready
for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell, have
been fixed; his suggestions have been incorporated with in the PEM production
code. Both updated HEPS and MEPS calibration data, along with the Level 3
code changes, will influence the in-situ energy deposition predictions, so new
versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues with no
problems. The MPG continues to assist the Project Scientistwith the scheduling
of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
2 days SEN (5.0) PROD (1992) 2R-2C=0L
Data received and cataloged at the CDHF in the month of March 2000 has been
previewed for HALOE, NMC, and UKMO in preparation for the DAAC.
CDHF REHOST:
- Successful processing on Rehost System:
Attitude Extraction, Observatory Log, Sensor Extraction, HALOE, MLS,
SUSIM, Solstice, and WINDII.
Preparing to test HRDI, ISAMS V12.0 and PEM V3.32 software on the
Rehost System.
***********
FOT
***********
REPORTING PERIOD
This report covers 13 May 2000 (Orbit 47388, GMT Day 134) through 19 May 2000
(Orbit 47491, GMT Day 140).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
WINDII 140/0008 47477 On
The beta angle ranged between the angles of 69.3 degrees to 67.3 degrees for
this report period. The beta angle is now decreasing toward a minimum angle of
+0.0 degrees on 05 June 2000 (DOY 157).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.8 to 1.0 degrees C. Additional battery performance data for each
battery on 12 May 2000 and 19 May 2000 are provided in Appendix D.
The clock error ranged from -12.1 to 11.0 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% OUT-OF-LIMITS = 02
% ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% FOT = 05
% SUSIM = 04
% HRDI = 03
% WINDII = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 02
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
- Transponder A: 00/091 - 2287.496948
- Transponder B: 00/091 - 2287.499113
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/139 show the observatory data loss to be 295 hours, 38
minutes, 10.561 seconds (an increase of 0 minutes 0.00 seconds since the last
report period). This is a 0.40 percent data loss which equals a 99.60 percent
data capture for the mission.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT was finally able to load the new TMONs into the simulator this week.
Testing of these TMONs has already begun and will continue into next week.
Once the TMONs have been thoroughly tested, they will be presented for review
and approval before uplink.
The FOT continues to develop an online procedure for safehold should it happen
before the new TMONs are in place or during and OBC halt. The NASA power
engineers were able to calculate the proper offset angle for this procedure.
This was the last calculation to be made so the procedure can now be finalized.
MEETINGS
The FOT attended the UARS FSW Meeting on 15 May 2000. The main topic discussed
was the three TDRS patch to increase the data coverage for UARS. The FSW teams
continues to analyze the GRO patch to see if it can be adapted to UARS. They
also plan to talk with the CMS software team to develop an implementation plan.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for June 05, 2000 (DOY 157). This maneuver
will be a Forward to Reverse Flight maneuver (Northern Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 19 May 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
47399 134/1922 00-212 HRDI SAFEHOLD - HI BARREL TEMP
47418 136/0141 00-213 OBC Clk Adj fr 41.2 to 20.6 msec/day
47444 137/1846 00-214 OBC Clk Adj fr 20.6 to 41.2 msec/day
47444 137/1928 00-215 HRDI SAFEHOLD - HI BARREL TEMP
47462 139/0005 00-216 SUSIM EXCEPTION CLEAR
47466 139/0551 00-217 NBTR PLAYBACK
47470 139/1207 00-218 HRDI SAFEHOLD - HI BARREL TEMP
47472 139/1611 00-219 SUSIM SUN ACTIVATE CMD
47473 139/1752 00-220 SUSIM DET OFF
47477 140/0008 00-221 WINDII POWER ON
47479 140/0228 00-222 OBC Clk Adj fr 41.2 to 20.6 msec/day
47480 140/0412 00-223 SUSIM SUSIM SUN ACTIVATE
47481 140/0553 00-217 NBTR PLAYBACK
_______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 19 May 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPTSBARTMP R-LO HRDI BARREL TMP 47399,421,462-72,81-91
HRPTSBARTMP Y-LO HRDI BARREL TMP 47405-08,417-20,432-39,461
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
-None this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 19 May 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
47418 136/01:42 (15 May 00) from 41.6 to 20.6 msec/day
47444 137/18:51 (16 May 00) from 20.6 to 41.6 msec/day
47479 140/02:31 (19 May 00) from 41.6 to 20.6 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 19 May 2000
12 May 2000 - GMT Day 133 Beta = 69.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- ---------- ---------- ------- ------- ------
1 +0.0/-22.4 -2.73/-4.31 +0.0/+0.0 15.8 0.000 0.0
2 +16.8/-11.2 +3.23/+1.97 +18.4/-14.0 27.0 1.040 15.0
3 +16.8/-22.4 +2.60/+1.34 +17.6/-13.6 27.0 1.037 14.8
19 May 2000 - GMT Day 140 Beta = 67.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- --------- ---------- ------- ------- ------
1 +11.2/-5.6 -3.68/-4.62 +0.0/+0.0 15.8 0.000 0.0
2 +16.8/-11.2 +3.87/+2.92 +23.2/-14.4 27.5 1.070 16.5
3 +16.8/-22.4 +2.92/+1.97 +22.4/-13.6 27.5 1.068 16.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 19 May 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-021 FOT 47408 135/1005 14170 GENERIC LATE ACQ
00-021 FOT 47458 138/1707 14170 GENERIC LATE ACQ
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiv 512
00-007 FOT see list May 00 14170 Generic Late Acq.
SEVNTFW00/20